# Systematically Measuring Ultra-Diffuse Galaxies. VIII. Misfits, Miscasts, and Miscreants

### v1

@@ 1 Introduction
In our own line of investigation (Zaritsky et al., 2019, 2022, 2023a), where we focus on compiling and then analyzing a large sample of physically large, low surface brightness galaxies (otherwise known as ultra-diffuse galaxies or UDGs; van Dokkum et al., 2015), we have faced such classification challenges and, on occasion, been surprised (Zaritsky et al., 2023b).
@@ 2 Data
The primary data source for the SMUDGes survey is the southern and northern sky images of the DESI Legacy Survey (Dey et al., 2019). We described the procedure that we use to identify the low-surface sources that comprise the SMUDGes catalog in a series of papers (Zaritsky et al., 2019, 2021, 2022, 2023a). As part of Zaritsky et al. (2023a), in which we present our completed catalog, we visually examined each of the 7070 candidates and flagged those that we found to be the most egregious misclassifications. Here, we revisit the entire list of candidates, appraise our classification of all, and categorize the objects that we identify as problematic. We now go beyond a simple binary flag and provide our best estimate of what the problematic detections might actually be. We will discuss the more interesting categories of SMUDGes, including mergers of low surface brightness galaxies, dwarf ring galaxies, and tidally distorted systems.
@@ 3.1.1 The Tidally Distorted
In this class we place those galaxies that appear to have been galaxies prior to the recent, ongoing interaction but whose current appearance we interpret to be under the influence of tidal forces. In some cases, such as those in Figure 1 where there is a clear nucleus and the shape is consistent with the presence of tidal tails, the assignment is relatively straightforward. Not all cases are so clear. If there is no nucleus and if the galaxy is closer to the larger host galaxy, then it can be quite difficult to differentiate this class of galaxy from the next class that we will be discussing. We classify 57 SMUDGes as tidally distorted and set them to have Class $=$ TD. All of our classification labels are summarized in Table 2.
@@
Background (generic)B
Background binaryBB
Background groupBG
Background tidalBT
CirrusC
DuplicateDU
Emission line regionE
UDG candidateG
Outer diskOD
Ongoing mergerOM
Post mergerPM
Ring galaxyR
Star S
Smoke ring galaxySR
Tidal debrisT
Tidal distortionD
Tidal dwarfTDW
@@
Figure 1: Two examples of tidally distorted SMUDGes sources (Class = TD). Images drawn from the Legacy Survey. The galaxy name and the size of the image in arcsec is given in each frame.
@@ 3.1.2 The Tidally Induced
We classify 11 SMUDGes as tidal dwarf galaxies and assign the Class $=$ TDW. This number represents less than 0.14% of the overall catalog, which may suggest that the tidal dwarf UDG population is negligible. However, as already mentioned, these ten are only those with a clear tidal connection to the parent galaxy. Others may have a connection that is not bright enough for us to detect with our current images. Furthermore, those tidal UDGs that have survived for a sufficient time to move away from the vicinity of associated tidal debris will go undetected under our classification criteria. Finally, it is also possible that some systems with strong tidal features were rejected by our pipeline and excluded from the catalog. We conclude that there is yet little evidence to suggest that this is a common formation pathway, but we cannot reach a conclusion on what fraction of the overall population they comprise. None of the 11 we identify has a spectroscopic redshift measurement, but spectroscopic redshifts are available for nine of the ten parent galaxies. In all but two of the cases, the inferred $r_{e}$ is $>$ 7 kpc. Because these values are unusually large even for UDGs, the results suggest that the majority of these are not true galaxies. If these turn out to be gravitationally bound galaxies, then their measured $r_{e}$ values are likely to be highly inflated due to their current appearance.
@@ 3.2.1 Dwarf Ring Galaxies
The SMUDGes pipeline usually identified the localized overdensity within the ring as the candidate UDG. The cataloged coordinates are therefore generally not at what one might more convincingly label the galaxy center. This is evident from the apparent miscentering of some objects in Figure 5. As such, the derived Sérsic parameters are more likely to reflect the size and structure of the overdensity rather than of the entire galaxy, again requiring caution when using the cataloged parameters for these sources. Even so, for the seven galaxies with existing spectroscopic measurements, six have an inferred $r_{e}>1.5$ kpc.
@@ 3.4 Background Sources
Among SMUDGes we reject are those that are galaxies, or sometimes multiple galaxies, that we classify as being well in the background and therefore not true low surface brightness dwarfs. We define several subclasses of these including a general background classification, Class $=$ B, binary galaxies that have been merged into a single detection, Class $=$ BB, tidal interactions among distant galaxies, Class $=$ BT, and groups of distant galaxies, Class $=$ BG. We classify 131 sources into one of these background categories. Additional background sources may fall into some of the classes discussed below. Unfortunately, only seven among these have a measured spectroscopic redshift but their values do confirm that they lie significantly farther than the bulk of our objects (these seven all have $cz>16,000$ km s-1, with an average value of $\sim$ 61,000 km s-1).
@@ 3.4.1 Tidal Detritus
Sometimes there is a fine line between a tidal dwarf (§3.1.2) and an overdensity of loose tidal material seen in projection as a concentration. We classify 157 SMUDGes as tidal debris, class $=$ T. If we include the background galaxies with tidal features that we classify as Class $=$ BT, then the number increases to 193.
@@
Figure 7: Two examples of tidal streamers (Class = TS). These vary in morphology but generally consist of a filamentary distribution of distinct blue pockets of star formation. The example on the right is the most extreme case, while the one on the left is more typical. Images drawn from the Legacy Survey. The galaxy name and the size of the image in arcsec is given in each frame.
@@ 3.5 Remainders
In this final category, we combine several classes, all of which should be rejected from further consideration as UDG candidates. First, there are 8 SMUDGes that appear to be artifacts, such as diffraction spikes or reflections. These are set to Class $=$ A. Second, there are 7 SMUDGes that are overdensities within the outskirts of luminous, nearby star forming galaxies. We set Class $=$ OD for these outer disk sources. Third, there are 8 SMUDGes that are bright stars, and we set Class $=$ S for these. Fourth, there are 2 SMUDGes that are Galactic emission line regions along jets. Both of these sources are part of previously known sources. The first (SMDG0333056-255813) is part of NGC 1360, a planetary nebula (Goldman et al., 2004). The second (SMDG2322299-021040) is associated with a Herbig-Haro object (Zuckerman et al., 2008). We set Class $=$ E for these objects. Finally, we set one duplicate detection to Class $=$ DU (the other detection is retained as a viable UDG candidate, Class = G).
@@ 4 Discussion
Figure 8: Angular correlation function for SMUDGes candidates. We select only those objects of class = G and exclude those within the Coma, Fornax, and Virgo clusters as designated in Zaritsky et al. (2023a). There is no correction for masked survey regions.

### v2

@@ 1 Introduction
In our own line of investigation (Zaritsky et al., 2019, 2022, 2023a), where we focus on compiling and then analyzing a large sample of physically large, low surface brightness galaxies (otherwise known as ultra-diffuse galaxies or UDGs; van Dokkum et al., 2015), we have faced such classification challenges and, on occasion, been surprised (Zaritsky et al., 2023b). In Zaritsky et al. (2023a) we discuss the full SMUDGes sample, including the distribution in distance and environment, when those measurements are available. We are in the process of improving our measurement of environment and preparing additional measurements, such as a quantification of the recent star formation histories, which we will presented in upcoming work.
@@ 2 Data
The primary data source for the SMUDGes survey is the southern and northern sky images of the DESI Legacy Survey (Dey et al., 2019). We described the procedure that we use to identify the low-surface sources that comprise the SMUDGes catalog in a series of papers (Zaritsky et al., 2019, 2021, 2022, 2023a). As part of Zaritsky et al. (2023a), in which we present our completed catalog, we visually examined each of the 7070 candidates and flagged those that we found to be the most egregious misclassifications. Here, we review the entire list of candidates, evaluate our classification of all, and categorize the objects we identify as problematic. We now go beyond a simple binary flag and provide our best estimate of what the problematic detections might actually be. We will discuss the more interesting categories of SMUDGes, including mergers of low surface brightness galaxies, dwarf ring galaxies, and tidally distorted systems.
@@ 3.1.1 The Tidally Distorted
In this class we place those galaxies that appear to have been galaxies prior to the recent, ongoing interaction but whose current appearance we interpret to be under the influence of tidal forces. In some cases, such as those in Figure 1 where there is a clear nucleus and the shape is consistent with the presence of tidal tails, the assignment is relatively straightforward. Not all cases are so clear. If there is no nucleus and if the galaxy is closer to the larger host galaxy, then it can be quite difficult to differentiate this class of galaxy from the next class that we will be discussing. We classify 56 SMUDGes as tidally distorted and set them to have Class $=$ D for distorted. All of our classification labels are summarized in Table 2.
@@
Background (generic)B0.61
Background binaryBB0.38
Background groupBG0.35
Background tidalBT0.52
CirrusC1.15
DuplicateDU0.01
Emission line regionE0.03
UDG candidateG92.69
Outer diskOD0.10
Ongoing mergerOM0.11
Post mergerPM0.10
Ring galaxyR0.41
Star S0.11
Smoke ring galaxySR0.03
Tidal debrisT2.22
Tidal distortionD0.79
Tidal dwarfTDW0.14
@@
Figure 1: Two examples of tidally distorted SMUDGes sources (Class = D). Images drawn from the Legacy Survey. The galaxy name and the size of the image in arcsec is given in each frame.
@@ 3.1.2 The Tidally Induced
We classify 10 SMUDGes as tidal dwarf galaxies and assign the Class $=$ TDW. This number represents $\sim$ 0.14% of the overall catalog, which may suggest that the tidal dwarf UDG population is negligible. However, as already mentioned, these ten are only those with a clear tidal connection to the parent galaxy. Others may have a connection that is not bright enough for us to detect with our current images. Furthermore, those tidal UDGs that have survived for a sufficient time to move away from the vicinity of associated tidal debris will go undetected under our classification criteria, although one might expect many to fall back toward their parent and be either accreted or tidally disrupted. Finally, it is also possible that some systems with strong tidal features were rejected by our pipeline and excluded from the catalog. We conclude that there is yet little evidence to suggest that this is a common formation pathway, but we cannot reach a conclusion on what fraction of the overall population they comprise. None of the 10 we identify has a spectroscopic redshift measurement, but spectroscopic redshifts are available for nine of the ten parent galaxies. In all but two of the cases, the inferred $r_{e}$ is $>$ 7 kpc. Because these values are unusually large even for UDGs, the results suggest that the majority of these are not true galaxies. If these turn out to be gravitationally bound galaxies, then their measured $r_{e}$ values are likely to be highly inflated due to their current appearance.
@@ 3.2.1 Dwarf Ring Galaxies
The SMUDGes pipeline usually identified the localized overdensity within the ring as the candidate UDG. The cataloged coordinates are therefore generally not at what one might more convincingly label the galaxy center. This is evident from the apparent miscentering of some objects in Figure 5. As such, the derived Sérsic parameters are more likely to reflect the size and structure of the overdensity rather than of the entire galaxy, again requiring caution when using the cataloged parameters for these sources. Even so, for the seven galaxies with existing spectroscopic measurements, six have an inferred $r_{e}>1.5$ kpc. The surface brightness is ill-defined for such irregular galaxies and the completeness estimates presented in the catalog, based on the injection of smooth objects into the survey data, are likely incorrect as well.
@@ 3.4 Background Sources
Among SMUDGes we reject are those that are galaxies, or sometimes multiple galaxies, that we classify as being well in the background and therefore not true low surface brightness dwarfs. We define several subclasses of these including a general background classification, Class $=$ B, binary galaxies that have been merged into a single detection, Class $=$ BB, tidal interactions among distant galaxies, Class $=$ BT, and groups of distant galaxies, Class $=$ BG. We classify 132 sources into one of these background categories. Additional background sources may fall into some of the classes discussed below. Unfortunately, only seven among these have a measured spectroscopic redshift but their values do confirm that they lie significantly farther than the bulk of our objects (these seven all have $cz>16,000$ km s-1, with an average value of $\sim$ 61,000 km s-1).
@@ 3.4.1 Tidal Detritus
Sometimes there is a fine line between a tidal dwarf (§3.1.2) and an overdensity of loose tidal material seen in projection as a concentration. We classify 157 SMUDGes as tidal debris, class $=$ T. If we include the background galaxies with tidal features that we classify as Class $=$ BT, then the number increases to 194.
@@
Figure 7: All identified tidal streamers (Class = TS). These vary in morphology but generally consist of a filamentary distribution of distinct blue pockets of star formation. Images drawn from the Legacy Survey. The galaxy name and the size of the image in arcsec is given in each frame.
@@ 3.5 Remainders
In this final category, we combine several classes, all of which should be rejected from further consideration as UDG candidates. First, there are 8 SMUDGes that appear to be artifacts, such as diffraction spikes or reflections. These are set to Class $=$ A, for artifact. Second, there are 7 SMUDGes that are overdensities within the outskirts of luminous, nearby star forming galaxies. We set Class $=$ OD for these outer disk sources. Third, there are 8 SMUDGes that are bright stars, and we set Class $=$ S for these. Fourth, there are 2 SMUDGes that are Galactic emission line regions along jets. Both of these sources are part of previously known sources. The first (SMDG0333056-255813) is part of NGC 1360, a planetary nebula (Goldman et al., 2004). The second (SMDG2322299-021040) is associated with a Herbig-Haro object (Zuckerman et al., 2008). We set Class $=$ E for these objects. Finally, we set one duplicate detection to Class $=$ DU (the other detection is retained as a viable UDG candidate, Class = G).
@@ 4 Discussion
Figure 8: Angular correlation function for SMUDGes candidates, which represents the excess fractional probability over random of finding a projected companion as a function of the pair separation. We select only those objects of class = G and exclude those within the Coma, Fornax, and Virgo clusters as designated in Zaritsky et al. (2023a). There is no correction for masked survey regions. The rise toward small angular separations demonstrates an increasing statistical excess of close projected companions.